A simple random sample of 100 postal employees is used to test if the average time postal employees have worked for the postal service has changed from the value of 7.5 years recorded 20 years ago. The sample mean was download?verifier=QIf83KKkrICt6hbxAvfkRy = 7 years with a standard deviation of s = 2 years. Assume the distribution of the time the employees have worked for the postal service is approximately Normal. The hypotheses being tested are H0: μ = 7.5 versus Ha: μ ≠ 7.5. A one-sample t test will be used. What is a 95% confidence interval for μ, the population mean time the postal service employees have spent with the postal service?

Group of answer choices

7 ± 2

7 ± 1.984

7 ± 0.4

7 ± 0.2
Answer #1

solutions - Given that, x = s=2 n=100 df= n-1 =100-1= 99 At a5 olo confidence level the tis L=1-95 % L=1-0.95 = 0.05 20.05 0.7 10.4 option (c) is correct Ho:4= 7.5 Ha: el # 7.5 This is two-tailed test Tast statistic, to ensure t = 7- 7.5 2/ Vioo t= -